2) Calculate the depth in sea water where the pressure is 10 atm.
Answer #1

2) answer) here we have

Ptotal=Po+\rhogh

Po=atmospheric pressure=1.013*105Pa

Ptotal=10*1.013*105pa

density of seawater=1030kg/m3

10*1.013*105pa=1.013*105pa+1030*9.8*h

9.117*105pa=10094*h

h=911700/10094=90.3 m

so the answer is 90.3 m